please answer all the following quesitons and show all working
1. Find the general solution to the following homogeneous differential equations. y - 2y + y = 0 (b) 9y + 6y + y = 0 4y
Answer #1

a co Sul yl_qyl +y= The Auxillay Eqnod Of is m²_qm+1=0 Im-lito m=I, Hemes general soln of DE is given by y=(4++)et qy+ 691+y© 4y+124 +9y=0 The Auxilley Eq is 4m2+ 12m +9=0 (2m +3 ]²=0 = m = -3% -3/2 so y = (4+Gt) este d Y_64+96=0 6 m² 6m+g=0 Th